OpenStudy (anonymous):

ho to differentiate y= (lnx)^5 ho to differentiate y= (lnx)^5 @Mathematics

OpenStudy (asnaseer):

you can do it by parts, first ley w=lnx so you get:\[y=w^5\]then use the rule:\[dy/dx=dy/dw*dw/dx\]you should be able to solve this now

OpenStudy (anonymous):

Use the power rule: y = (lnx)^5 let u = lnx du/dx = 1/x y = u^5 dy/du = 5u^4 du/dx dy/dx = (5/x)(lnx)^4

OpenStudy (anonymous):

thanks....